OpenStudy (anonymous):

Solve the following application problem. The formula for the n th term b of a geometric series is b = arn-1. Find n when b = 1,024 a = 16, and r = 2.

OpenStudy (anonymous):

That's easy just plugh 1,024 instead of b and 16 instead of a and 2 instead of r and solve and get n :)

OpenStudy (anonymous):

Plug*

OpenStudy (anonymous):

I did that and got 3, and it`s wrong :(

OpenStudy (anonymous):

1024=16*2^(n-1) (divide by 16) 64=2^(n-1) (65=2^6) 2^6=2^(n-1) 6=n-1 , n=7

OpenStudy (anonymous):